Given the sides of a triangle x+y-6=0; 3x-5y+14=0; 5x-3y-14=0 Express the hieght of this triangle.
Answer by Alan3354(69443)

Given the sides of a triangle x+y-6=0; 3x-5y+14=0; 5x-3y-14=0 Express the hieght of this triangle.
---------
The sides can't be equations.
And, it says the 3 sides = 0.
So the height is zero.